OpenStudy (gabbi):

what is 60/100 in lowest terms

OpenStudy (anonymous):

0.6?

OpenStudy (anonymous):

0.6 or 3/5

OpenStudy (allank):

Since they've said term(s), 3/5 sounds okay.

OpenStudy (anonymous):

60/100 = 6/10=3/5